OpenStudy (anonymous):

The lunch platter at one restaurant costs $7. Write an equation that shows the total cost, c, of p people buying the lunch platter.

OpenStudy (misty1212):

one person, the cost is $7 two people it is \(2\times\$7=\$14\) three people it is \(3\times \$7=\$21\) four people it is \(4\times \$7=\$28\)

OpenStudy (misty1212):

for \(p\) people it is \(p\times \$7\) or \(7\times p\)

OpenStudy (misty1212):

most people would just write \[C=7p\]

OpenStudy (anonymous):

thank youu your so good at this omg!

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]

OpenStudy (anonymous):

byee

OpenStudy (misty1212):

ttyl